# Rethinking the Role of Infrastructure in Collaborative Perception

### v1

@@ Abstract
Collaborative Perception (CP) is a process in which an ego agent receives and fuses sensor information from surrounding vehicles and infrastructure to enhance its perception capability. To evaluate the need for infrastructure equipped with sensors, extensive and quantitative analysis of the role of infrastructure data in CP is crucial, yet remains underexplored. To address this gap, we first quantitatively assess the importance of infrastructure data in existing vehicle-centric CP, where the ego agent is a vehicle. Furthermore, we compare vehicle-centric CP with infra-centric CP, where the ego agent is now the infrastructure, to evaluate the effectiveness of each approach. Our results demonstrate that incorporating infrastructure data improves 3D detection accuracy by up to 10.87%, and infra-centric CP shows enhanced noise robustness and increases accuracy by up to 42.53% compared with vehicle-centric CP.
@@ 1 Introduction
The paper is structured as follows. In Section 2, we summarize previous studies of vehicle-centric and infra-centric CP. Section 3 overviews the process of CP, and Section 4 presents the experiment details and results. We conclude this paper in Section 5.
@@ 4.1 Experiment Setup
We modify the CP models to convert vehicle-centric CP to infra-centric CP. For all the dataset scenarios, we set the infrastructure as the ego agent and change the the detection range of $z$ to consider the height difference between the vehicle and the infrastructure. In V2XSet-I, we convert the range of $z$ from $z\in[-3,1]$ to $z\in[-5,-1]$, and, in V2X-Sim, we convert from $z\in[-3,2]$ to $z\in[-8.5,-3.5]$. We adjust the detection range of $x$ and $y$ of the ego agent. In vehicle-centric CP, we use the default detection range given in [26] for V2XSet and [12] for V2X-Sim. In detail, the detection range of V2XSet-W and V2XSet-I is $x\in[-140.8,140.8]$ and $y\in[-38.4,38.4]$, and the detection range of V2X-Sim is $x\in[-32,32]$ and $y\in[-32,32]$. In infra-centric CP, we only change the detection range in V2XSet-I. We use the square-shaped detection range of $x\in[-76.8,76.8]$ and $y\in[-76.8,76.8]$. The reason for this selection is explained in Section 4.3.1. In V2X-Sim, we set the detection range for infra-centric CP as the same as vehicle-centric CP because the default detection range is already square-shaped.
@@ 4.2 Study of Infrastructure Data in Vehicle-centric CP
V2X-ViT [26] 0.685 0.751 0.599 0.636 0.806 0.828 0.594 0.577
Where2comm [9] 0.715 0.793 0.633 0.662 0.770 0.801 0.577 0.573
ParCon [2] 0.746 0.817 0.665 0.688 0.781 0.809 0.630 0.618
@@
The results of the accuracy comparison between V2V CP and V2X CP are in Table 1. Regarding the detection accuracy on V2XSet, for all the models, V2X CP shows better accuracy than V2V CP. In the perfect setting, AP@0.7 increases by between 9.51% and 10.87%, and in the simple noise setting, AP@0.7 increases by between 3.49% and 6.11%. Regarding the detection accuracy on V2X-Sim, for all the models, V2X CP shows better accuracy in the perfect setting, enhanced by between 2.67% and 4.03%. However, in the simple noise setting, V2X CP has lower accuracy than V2V CP, decreased by between 0.70% and 2.89%.
In the simple noise setting, the accuracy of V2X CP is better in V2XSet and worse in V2X-Sim than that of V2V CP. The main cause is noise itself. Noise tends to be more severe when sensor data points are far from the sensor (e.g., for the same level of heading angle noise, data points farther from the sensor represent excessive movement compared to those closer to the sensor). For this reason, the infrastructure’s property of a broader sensor range makes the infrastructure more vulnerable to noise than the vehicle. Furthermore, noise makes perception ability more vulnerable with the smaller detection range. As mentioned in Section 4.1.3, V2X-Sim uses a smaller detection range than V2XSet. The smaller detection range is less likely to include the auxiliary agents’ data. This indicates that the infrastructure might rarely influence the ego vehicle to enhance perception, or infrastructure data with noise deteriorates the perception ability, which acts as an obstruction.
@@
Model CP Scene #1 Scene #3 Scene #4 Scene #7 Scene #9 Scene #11
V2X-ViT[26] V2V 0.920 0.866 0.561 0.774 0.663 0.781
V2X 0.841 0.782 0.762 0.932 0.744 0.749
Difference 0.079$\downarrow$ 0.084$\downarrow$ 0.201$\uparrow$ 0.158$\uparrow$ 0.081$\uparrow$ 0.032$\downarrow$
Where2comm[9] V2V 0.951 0.877 0.569 0.856 0.701 0.805
V2X 0.934 0.783 0.801 0.949 0.795 0.795
Difference 0.017$\downarrow$ 0.094$\downarrow$ 0.232$\uparrow$ 0.093$\uparrow$ 0.094$\uparrow$ 0.009$\downarrow$
ParCon[2] V2V 0.953 0.907 0.656 0.776 0.706 0.864
V2X 0.957 0.874 0.838 0.978 0.804 0.796
Difference 0.004$\uparrow$ 0.033$\downarrow$ 0.182$\uparrow$ 0.201$\uparrow$ 0.098$\uparrow$ 0.068$\downarrow$
@@
We compare V2V CP and V2X CP in the same specific scenarios on the V2XSet dataset. As shown in Table 2, the performance of V2V CP and V2X CP largely depends on scenarios. To analyze when V2X CP is more effective than V2V CP, we select two scenarios, Scene #4 and Scene #3, which are the most and least effective cases of V2X CP, respectively.
@@ 4.3 Study of Infra-centric CP
V2X-ViT [26] 0.909/0.851 0.895/0.793
Where2comm [9] 0.928/0.895 0.927/0.880
ParCon[2] 0.949/0.931 0.938/0.878
@@
V2X-ViT [26] 0.875/0.832 0.924/0.884
Where2comm [9] 0.857/0.789 0.940/0.895
ParCon[2] 0.897/0.853 0.956/0.933
@@
Depending on the type of the ego agent, the detection range may have to change. As shown in Tables 4 and 4, the rectangle shape of the detection range shows better performance than the square shape in V2X CP. On the other hand, the square shape performs better than the rectangle shape in I2X CP. In Figure 5 (a), the detection range for the ego vehicle in V2X CP should adopt a rectangular shape, aligning with the road’s layout and the vehicle’s forward movement. When the square-shaped detection range is applied, the heading direction length is shorter than the rectangle-shaped detection range, thereby missing the objects far away from the ego vehicle along the heading direction (Boundary 1 in Fig 5 (a)). Also, the square-shaped detection range makes V2X CP detect unnecessary objects that are unlikely to interact with the ego vehicle (Boundary 2 in Fig 5 (a)). Conversely, in Figure  5 (b), the detection range of the ego infrastructure in I2X CP should be square-shaped. This range leverages the property that infrastructure sensors are elevated and can monitor all directions equally. When the rectangle-shaped detection range is applied, it includes an area that is not of interest (Boundary 3 in Fig 5 (b)). Also, the rectangle-shaped detection range makes I2X CP unable to fully observe the responsible intersection area (Boundary 4 in Fig 5 (b)).
@@
To identify the effects of the type of an ego agent, we compare the detection accuracy of V2X CP and I2X CP as shown in Table 5. In the perfect setting on the V2XSet, all the models show that I2X CP outperforms V2X CP in accuracy. Also, I2X CP is more effective in the simple noise setting, showing a rate of increase from 2.50% to 7.55% on V2XSet and from 29.75% to 42.53% on V2X-Sim. Unlike the comparison between V2V and V2X CP in section 4.2.2, it is noteworthy that I2X CP always exhibits better performance in the simple noise setting. The finding indicates that the ego infrastructure has noise robustness. The standalone infrastructure performs well, and based on this infrastructure’s ability, I2X CP maintains its performance even in the noise. Regarding the comparison between datasets, the change rate in V2X-Sim is significantly larger than that in V2XSet. As we mentioned in Sec. 4.1.3, the phenomenon might stem from the detection range of V2X-Sim, which is almost square-shaped and corresponds with a region at an intersection. Thus, I2X CP can cover almost every region within the detection range by minimizing occlusion. This supports our logic in shaping the detection range in section 4.3.1.
@@
No Fusion 0.675 0.718 0.675 0.718 0.517 0.795 0.517 0.795
V2X-ViT [26] 0.851 0.884 0.803 0.860 0.828 0.873 0.577 0.822
Where2comm [9] 0.895 0.895 0.815 0.876 0.801 0.860 0.573 0.814
ParCon [2] 0.931 0.933 0.848 0.870 0.809 0.834 0.618 0.802
@@
Based on the model trained with the harsh noise setting, which is detailed in [2], we compare the noise sensitivity of V2X CP and I2X CP. As shown in Fig. 6, all the I2X CP models show better accuracy than the V2X CP models in all of the noises on the V2Xset dataset. Also, the decline rate of the accuracy of I2X CP is lower than that of V2X CP as the noise worsens. Likewise, all the I2X CP models outperform the V2X CP models and have a lower decline rate in V2X-Sim, as shown in Fig. 7. In addition, the results on V2X-Sim show a significantly large accuracy difference between I2X CP and V2X CP, even in the zero noise. These results indicate that changing the ego agent to infrastructure allows for improved noise robustness because the data of an ego agent is not affected by communication noises; the ego infrastructure utilizes noise-free data itself, thereby enabling I2X CP to show good performance in noisy environments and maintain performance even when the received data worsens.
@@
As shown in Table 6, the accuracy of I2X CP is different depending on the scenarios, and the degree of effectiveness becomes more obvious than Table 2, which is accuracy comparison between V2V CP and V2X CP. To analyze when I2X CP outperforms V2X CP, we choose two scenarios: Scene #2 and Scene #12, which are the most and least effective cases of I2X CP, respectively.
@@
V2X-ViT[26] V2X 0.637 0.575 0.640 0.729 0.634 0.579
I2X 0.982 0.125 0.781 0.424 0.824 0.061
Difference 0.345$\uparrow$ 0.451$\downarrow$ 0.141$\uparrow$ 0.304$\downarrow$ 0.190$\uparrow$ 0.518$\downarrow$
Where2comm[9] V2X 0.684 0.641 0.706 0.835 0.708 0.713
I2X 0.983 0.173 0.842 0.446 0.849 0.123
Difference 0.299$\uparrow$ 0.468$\downarrow$ 0.136$\uparrow$ 0.389$\downarrow$ 0.141$\uparrow$ 0.589$\downarrow$
ParCon[2] V2X 0.750 0.696 0.695 0.859 0.729 0.692
I2X 0.975 0.239 0.843 0.513 0.870 0.117
Difference 0.225$\uparrow$ 0.457$\downarrow$ 0.148$\uparrow$ 0.346$\downarrow$ 0.140$\uparrow$ 0.575$\downarrow$
@@
Figure 8: Effective case of I2X CP (Scene #2). The point clouds from the vehicle’s LiDAR are indicated as white dots, and the point clouds from the infrastructure’s LiDAR are indicated as yellow dots. The green bounding boxes are ground truth objects, and the red bounding boxes are predicted objects. Both boxes are located in the ego agent’s detection range. E means the ego agent and A means aux agents.

### v2

@@ Abstract
Collaborative Perception (CP) is a process in which an ego agent receives and fuses sensor information from surrounding vehicles and infrastructure to enhance its perception capability. To evaluate the need for infrastructure equipped with sensors, extensive and quantitative analysis of the role of infrastructure data in CP is crucial, yet remains underexplored. To address this gap, we first quantitatively assess the importance of infrastructure data in existing vehicle-centric CP, where the ego agent is a vehicle. Furthermore, we compare vehicle-centric CP with infra-centric CP, where the ego agent is now the infrastructure, to evaluate the effectiveness of each approach. Our results demonstrate that incorporating infrastructure data improves 3D detection accuracy by up to 10.30%, and infra-centric CP shows enhanced noise robustness and increases accuracy by up to 46.47% compared with vehicle-centric CP.
@@ 1 Introduction
The paper is structured as follows. In Sec. 2, we summarize previous studies of vehicle-centric and infra-centric CP. Sec. 3 overviews the process of CP, and Sec. 4 presents the experiment details and results. We conclude this paper in Sec. 5.
@@ 4.1 Experiment Setup
We modify the CP models to convert vehicle-centric CP to infra-centric CP. For all the dataset scenarios, we set the infrastructure as the ego agent and change the the detection range of $z$ to consider the height difference between the vehicle and the infrastructure. In V2XSet-I, we convert the range of $z$ from $z\in[-3,1]$ to $z\in[-5,-1]$, and, in V2X-Sim, we convert from $z\in[-3,2]$ to $z\in[-8.5,-3.5]$. We adjust the detection range of $x$ and $y$ of the ego agent. In vehicle-centric CP, we use the default detection range given in [26] for V2XSet and [12] for V2X-Sim. In detail, the detection range of V2XSet-W and V2XSet-I is $x\in[-140.8,140.8]$ and $y\in[-38.4,38.4]$, and the detection range of V2X-Sim is $x\in[-32,32]$ and $y\in[-32,32]$. In infra-centric CP, we only change the detection range in V2XSet-I. We use the square-shaped detection range of $x\in[-76.8,76.8]$ and $y\in[-76.8,76.8]$. The reason for this selection is explained in Sec. 4.3.1. In V2X-Sim, we set the detection range for infra-centric CP as the same as vehicle-centric CP because the default detection range is already square-shaped.
@@ 4.2 Study of Infrastructure Data in Vehicle-centric CP
V2X-ViT [26] 0.673 0.722 0.548 0.564 0.745 0.775 0.578 0.594
Where2comm [9] 0.713 0.785 0.572 0.580 0.733 0.752 0.559 0.555
ParCon [2] 0.734 0.810 0.630 0.646 0.793 0.829 0.615 0.621
@@
The results of the accuracy comparison between V2V CP and V2X CP are in Tab. 1. Regarding the detection accuracy on V2XSet, for all the models, V2X CP shows better accuracy than V2V CP. In the perfect setting, AP@0.7 increases by between 7.29% and 10.30%, and in the simple noise setting, AP@0.7 increases by between 1.43% and 2.98%. Regarding the detection accuracy on V2X-Sim, for all the models, V2X CP shows better accuracy in the perfect setting, enhanced by between 2.53% and 4.55%, and in the simple noise setting, AP@0.7 of V2X-ViT and ParCon increase by 2.71% and 0.86%, respectively. However, Where2comm has lower accuracy about V2X CP than V2V CP, decreased by 0.73%.
In the simple noise setting, the accuracy of V2X CP is better in V2XSet and worse in V2X-Sim than that of V2V CP. The main cause is noise itself. Noise tends to be more severe when sensor data points are far from the sensor (e.g., for the same level of heading angle noise, data points farther from the sensor represent excessive movement compared to those closer to the sensor). For this reason, the infrastructure’s property of a broader sensor range makes the infrastructure more vulnerable to noise than the vehicle. Furthermore, noise makes perception ability more vulnerable with the smaller detection range. As mentioned in Sec. 4.1.3, V2X-Sim uses a smaller detection range than V2XSet. The smaller detection range is less likely to include the auxiliary agents’ data. This indicates that the infrastructure might rarely influence the ego vehicle to enhance perception, or infrastructure data with noise deteriorates the perception ability, which acts as an obstruction.
@@
Model CP Scene #1 Scene #3 Scene #4 Scene #8 Scene #9 Scene #11
V2X-ViT[26] V2V 0.925 0.852 0.579 0.634 0.668 0.768
V2X 0.907 0.742 0.776 0.698 0.724 0.746
Difference 0.018$\downarrow$ 0.110$\downarrow$ 0.197$\uparrow$ 0.064$\uparrow$ 0.056$\uparrow$ 0.022$\downarrow$
Where2comm[9] V2V 0.954 0.889 0.602 0.648 0.690 0.818
V2X 0.956 0.804 0.809 0.733 0.786 0.822
Difference 0.002$\uparrow$ 0.085$\downarrow$ 0.208$\uparrow$ 0.084$\uparrow$ 0.097$\uparrow$ 0.005$\uparrow$
ParCon[2] V2V 0.953 0.892 0.680 0.710 0.699 0.865
V2X 0.971 0.857 0.854 0.798 0.818 0.854
Difference 0.019$\uparrow$ 0.035$\downarrow$ 0.174$\uparrow$ 0.089$\uparrow$ 0.120$\uparrow$ 0.012$\downarrow$
@@
We compare V2V CP and V2X CP in the same specific scenarios on the V2XSet dataset. As shown in Tab. 2, the performance of V2V CP and V2X CP largely depends on scenarios. To analyze when V2X CP is more effective than V2V CP, we select two scenarios, Scene #4 and Scene #3, which are the most and least effective cases of V2X CP, respectively.
@@ 4.3 Study of Infra-centric CP
V2X-ViT [26] 0.926/0.875 0.899/0.815
Where2comm [9] 0.945/0.913 0.933/0.881
ParCon[2] 0.938/0.902 0.944/0.904
@@
V2X-ViT [26] 0.890/0.828 0.915/0.882
Where2comm [9] 0.882/0.800 0.947/0.910
ParCon[2] 0.885/0.818 0.936/0.908
@@
Depending on the type of the ego agent, the detection range may have to change. As shown in Tab. 4 and Tab. 4, the rectangle shape of the detection range shows better performance than the square shape in V2X CP. On the other hand, the square shape performs better than the rectangle shape in I2X CP. In Fig. 5 (a), the detection range for the ego vehicle in V2X CP should adopt a rectangular shape, aligning with the road’s layout and the vehicle’s forward movement. When the square-shaped detection range is applied, the heading direction length is shorter than the rectangle-shaped detection range, thereby missing the objects far away from the ego vehicle along the heading direction (Boundary 1 in Fig. 5 (a)). Also, the square-shaped detection range makes V2X CP detect unnecessary objects that are unlikely to interact with the ego vehicle (Boundary 2 in Fig. 5 (a)). Conversely, in Fig. 5 (b), the detection range of the ego infrastructure in I2X CP should be square-shaped. This range leverages the property that infrastructure sensors are elevated and can monitor all directions equally. When the rectangle-shaped detection range is applied, it includes an area that is not of interest (Boundary 3 in Fig. 5 (b)). Also, the rectangle-shaped detection range makes I2X CP unable to fully observe the responsible intersection area (Boundary 4 in Fig. 5 (b)).
@@
To identify the effects of the type of an ego agent, we compare the detection accuracy of V2X CP and I2X CP as shown in Tab. 5. In the perfect setting on the V2XSet, all the models show that I2X CP outperforms V2X CP in accuracy. Also, I2X CP is more effective in the simple noise setting, showing a rate of increase from 16.65% to 19.13% on V2XSet and from 37.51% to 46.47% on V2X-Sim. Unlike the comparison between V2V and V2X CP in sec. 4.2.2, it is noteworthy that I2X CP always exhibits better performance in the simple noise setting at AP@0.7. The finding indicates that the ego infrastructure has noise robustness. The standalone infrastructure performs well, and based on this infrastructure’s ability, I2X CP maintains its performance even in the noise. Regarding the comparison between datasets, the change rate in V2X-Sim is significantly larger than that in V2XSet. As we mentioned in Sec. 4.1.3, the phenomenon might stem from the detection range of V2X-Sim, which is almost square-shaped and corresponds with a region at an intersection. Thus, I2X CP can cover almost every region within the detection range by minimizing occlusion. This supports our logic in shaping the detection range in Sec. 4.3.1.
@@
No Fusion 0.602 0.718 0.602 0.718 0.517 0.795 0.517 0.795
V2X-ViT [26] 0.875 0.882 0.716 0.853 0.775 0.879 0.594 0.820
Where2comm [9] 0.913 0.910 0.724 0.855 0.752 0.868 0.555 0.813
ParCon [2] 0.902 0.908 0.737 0.859 0.829 0.896 0.621 0.853
@@
Based on the model trained with the harsh noise setting, we compare the noise sensitivity of V2X CP and I2X CP. As shown in Fig. 6, all the I2X CP models show better accuracy than the V2X CP models in all of the noises on the V2Xset dataset. Also, the decline rate of the accuracy of I2X CP is lower than that of V2X CP as the noise worsens. Likewise, all the I2X CP models outperform the V2X CP models and have a lower decline rate in V2X-Sim, as shown in Fig. 7. In addition, the results on V2X-Sim show a significantly large accuracy difference between I2X CP and V2X CP, even in the zero noise. These results indicate that changing the ego agent to infrastructure allows for improved noise robustness because the data of an ego agent is not affected by communication noises; the ego infrastructure utilizes noise-free data itself, thereby enabling I2X CP to show good performance in noisy environments and maintain performance even when the received data worsens.
@@
As shown in Tab. 6, the accuracy of I2X CP is different depending on the scenarios, and the degree of effectiveness becomes more obvious than Tab. 2, which is accuracy comparison between V2V CP and V2X CP. To analyze when I2X CP outperforms V2X CP, we choose two scenarios: Scene #2 and Scene #12, which are the most and least effective cases of I2X CP, respectively.
@@
V2X-ViT[26] V2X 0.813 0.824 0.876 0.765 0.728 0.733
I2X 0.990 0.476 0.913 0.997 0.134 0.069
Difference 0.176$\uparrow$ 0.347$\downarrow$ 0.037$\uparrow$ 0.233$\uparrow$ 0.595$\downarrow$ 0.664$\downarrow$
Where2comm[9] V2X 0.849 0.900 0.862 0.924 0.770 0.786
I2X 0.993 0.514 0.920 0.998 0.178 0.096
Difference 0.144$\uparrow$ 0.386$\downarrow$ 0.059$\uparrow$ 0.074$\uparrow$ 0.593$\downarrow$ 0.690$\downarrow$
ParCon[2] V2X 0.870 0.920 0.854 0.833 0.824 0.840
I2X 0.990 0.445 0.944 0.997 0.198 0.116
Difference 0.119$\uparrow$ 0.475$\downarrow$ 0.090$\uparrow$ 0.164$\uparrow$ 0.626$\downarrow$ 0.725$\downarrow$
@@
Figure 8: Effective case of I2X CP (Scene #7). The point clouds from the vehicle’s LiDAR are indicated as white dots, and the point clouds from the infrastructure’s LiDAR are indicated as yellow dots. The green bounding boxes are ground truth objects, and the red bounding boxes are predicted objects. Both boxes are located in the ego agent’s detection range. E means the ego agent and A means aux agents.